# What the flock knows that the birds do not: exploring the emergence of joint agency in multi-agent active inference

### v1

@@ A.4 Spectral Identification of Markov Blankets via Fiedler Vector Analysis
For community detection applications, we consider the Markov blanket separating two communities $C_{1}$ and $C_{2}$ as the subset of boundary nodes that minimizes information flow between the communities while maintaining graph connectivity. To find these boundary nodes, we determine the algebraic connectivity (also known as Fiedler eigenvalue) of the graph by computing the second-smallest eigenvalue of the Laplacian matrix of $F$ defined as L=DAL=D-A where $D=\text{diag}(\sum_{j}A_{ij})$ is the degree matrix. The Fiedler vector $y_{2}$ (fiedler1973algebraic) is the eigenvector corresponding to the second smallest eigenvalue $\lambda_{2}$ of $L$, such that $Ly_{2}=\lambda_{2}y_{2}$.
@@ Acknowledgments
All the authors contributed to the conceptualization and writing of the manuscript.

### v2

@@ A.4 Spectral Identification of Markov Blankets via Fiedler Vector Analysis
For community detection applications, we consider the Markov blanket separating two communities $C_{1}$ and $C_{2}$ as the subset of boundary nodes that minimizes information flow between the communities while maintaining graph connectivity. To find these boundary nodes, we determine the algebraic connectivity (also known as Fiedler eigenvalue) of the graph by computing the second-smallest eigenvalue of the Laplacian matrix of $F$ defined as L=DAL=D-A where $D=\text{diag}(\sum_{j}A_{ij})$ is the degree matrix. The Fiedler vector $y_{2}$ (Fiedler, 1973) is the eigenvector corresponding to the second smallest eigenvalue $\lambda_{2}$ of $L$, such that $Ly_{2}=\lambda_{2}y_{2}$.
@@
A.5 Information-Theoretic Analysis of Agent Interactions
Let $X_{i}^{k}(t)$ denote the state of agent $i$ at time $t$ in simulation $k$, where each simulation is a sample from the same stochastic process, used subsequently to estimate probabilities. The agent’s state is encoded as an integer between 0 and 3, corresponding to its heading direction. The predator position in simulation $k$ at time $t$ is denoted by $Y^{k}(t)$. This variable is discretized into four spatial quadrants (hence also taking values between 0 and 3) and, depending on the analysis, can refer to either the first or the second predator in the simulation.
For each pair of agents $i$ and $j$, we compute the univariate and multivariate mutual information, respectively $I_{i}(t)$ ($I_{j}(t)$) and $I_{ij}(t)$, between their states $X_{i}^{k}(t)$, $X_{j}^{k}(t)$ and the target variable $Y^{k}(t)$:
$I_{ij}(t)$ $=I\bigl(X_{i}(t),X_{j}(t);Y(t)\bigr),$ (11)
$I_{i}(t)$ $=I\bigl(X_{i}(t);Y(t)\bigr),$
$I_{j}(t)$ $=I\bigl(X_{j}(t);Y(t)\bigr),$
All mutual information terms are evaluated at each time step $t$, treating $X_{i}(t)$, $X_{j}(t)$, and $Y(t)$ as random variables whose realizations are drawn from the ensemble of simulations $\{(X_{i}^{k}(t),X_{j}^{k}(t),Y^{k})\}_{k}(t)$.
We then perform a standard partial information decomposition (PID) using the following expressions:
$R_{ij}(t)$ $=\min\bigl\{I_{i}(t),I_{j}(t)\bigr\},$ (12)
$U_{i}(t)$ $=I_{i}(t)-R_{ij}(t),$
$U_{j}(t)$ $=I_{j}(t)-R_{ij}(t),$
$S_{ij}(t)$ $=I_{ij}(t)-R_{ij}(t)-U_{i}(t)-U_{j}(t),$
where $R_{ij}$, $U_{i}$, $U_{j}$, and $S_{ij}$ represent the redundant — using the definition present in (Williams and Beer, 2010) —, unique (agent $i$), unique (agent $j$), and synergistic information components, respectively. These quantities are averaged across all agent pairs at a specified interaction distance; in the absence of a fixed distance, they are averaged over all possible pairs.
To assess the statistical significance of the obtained information components, we estimate null distributions by preserving the source variables $X_{i}$ and $X_{j}$ while applying a temporal shuffle to $Y(t)$. This procedure is repeated $N=500$ times, and the mutual information and PID atoms are recomputed for each surrogate series. This method generates a null distribution under the hypothesis of no systematic information sharing between agent states and the predator’s position, following the surrogate-data approach commonly used in information-theoretic connectivity analysis (Vicente et al., 2011).
@@ Acknowledgments
All authors contributed to the conceptualization and writing of the manuscript.
@@
Data availability
The source code to reproduce the experiments is available at https://github.com/dommai/flock_knows_what_birds_dont